a random sample of n=755 us cell phobe users age 18 and older in may 2011 found that the average number of text messages sent or received per day is 41.5 messages with standard error about 6.1

find 95% confidence interval for the mean number of text messages. the 95% confidence interval is?
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 41.5

standard error =    = 6.1

Sample size = n = 755

At 95% confidence level

= 1 - 95%  

= 1 - 0.95 =0.05

/2 = 0.025

Z/2 = Z0.025 = 1.96


Margin of error = E = Z/2 *

= 1.96 * 6.1

= 11.96

At 95% confidence interval estimate of the population mean is,

  ± E

41.5 ± 11.96   

( 29.54 , 53.46)